# Are all Binary Black Holes Detected from LIGO-Virgo-KAGRA Follow the Universal Time-Delay Distributions? Likely Not

### v1

@@ 1 Introduction
We apply this framework to the GWTC-4 (LIGO-VIRGO-KAGRA Collaboration, 2025b) catalog (including GWTC-3 (Abbott et al., 2023)), detected by the LIGO(Abbott et al., 2016b), Virgo(Acernese et al., 2015), and KAGRA(Akutsu et al., 2021) (LVK) collaboration. The analysis reveals a tentative evidence of mass dependence of the DTD: lower-mass binaries (-) favor DTDs that appear broadly flat in the space across 1-10 Gyr, consistent with mergers occurring over a wide range of evolutionary timescales. By contrast, higher-mass binaries (-) exhibit a tentative evidence of non-power law behavior with a sharply peaked distributions at short delays (3-5 Gyr), indicative of rapid post-formation mergers. These findings provide new insights into mass-dependent binary evolution and demonstrate the power of non-parametric approaches for extracting detailed astrophysical information from GW observations.
@@ 2.3 Inference of Local Merger Rate Density
Once the best-fitting DDT shape has been identified for each mass bin through the Bayesian framework described above, we proceed to infer the local merger rate density . For this inference, we note that the shape of determines the redshift evolution of the merger rate, while acts as a normalization parameter that scales the overall amplitude. With the best-fitting trajectory fixed, the inference of becomes a single-parameter problem for each mass bin. We adopt a flat prior on ranging from 1 to 100 Gpc-3 yr-1. The likelihood function follows the same Poisson formulation, with the expected counts now explicitly dependent on . This yields both point estimates and uncertainties for the local BBH merger rate density across different mass bins.
@@ 4 Results
In contrast, higher-mass binaries (-) exhibit peaked DTDs, with merger probabilities strongly concentrated at short to intermediate delays (3-5 Gyr) and a sharp suppression of long delays. The decay in the high for the high mass bin is coming from the fact that the merger rate is increasing as one moves from lower redshifts to higher redshift (see figure 2 red solid line.). Whereas at the lower delay time (for Gyrs), the constraints are weak (exhibiting a large error) due to non-detection of high redshift sources impacted by the selection function. In figure 2 we show the normalized redshift distributions of BBH mergers, comparing the observed GW catalog data (solid line) with the reconstructed trajectories (dashed line) derived from the best-evidence DTD inferred using our grid-based technique. The remarkable agreement between the two provides a strong self-consistency check on our framework, demonstrating that the inferred DTD not only yield statistically favored trajectories in space but also naturally reproduce the observed cosmic evolution of the merger population.

### v2

# Are all Binary Black Holes Detected by LIGO-Virgo-KAGRA Following the Universal Time-Delay Distributions? Probably Not

@@ 1 Introduction
We apply this framework to the GWTC-4 (LIGO-VIRGO-KAGRA Collaboration, 2025b) catalog (including GWTC-3 (Abbott et al., 2023)), detected by the LIGO(Abbott et al., 2016b), Virgo(Acernese et al., 2015), and KAGRA(Akutsu et al., 2021) (LVK) collaboration. The analysis reveals a tentative evidence of mass dependence of the DTD: lower-mass binaries (-) favor DTDs that appear broadly flat in the space across 1-10 Gyr, consistent with mergers occurring over a wide range of evolutionary timescales. By contrast, higher-mass binaries (-) exhibit a tentative evidence of non-power law behavior with a sharply peaked distributions at short delays (2-6 Gyr), indicative of rapid post-formation mergers. These findings provide new insights into mass-dependent binary evolution and demonstrate the power of non-parametric approaches for extracting detailed astrophysical information from GW observations.
@@ 2 Theoretical Framework and Grid Construction
It is important to note that in this work we assume the merger rate follows the cosmic star formation history convolved with a DTD. However, since the birth metallicity distribution evolves with redshift and the formation of massive BBHs is strongly suppressed at high metallicities (Belczynski et al., 2010; Mapelli, 2016; Lamberts, 2018; Mukherjee, 2022; Afroz & Mukherjee, 2025a; Neijssel et al., 2019), this simplification does not explicitly account for metallicity-dependent formation efficiency. Our main goal in this paper is to identify whether there is any deviation from an universal power-law form of the delay-time distribution in a mass-dependent way. Moreover, it is also important to note that the metallicity-dependent star formation rate as a function of redshift is not well measured from observations. As a result, in this analysis which is the first-of-its-kind from GWTC-4, we do not allow a metallicity dependent star formation rate. Given the limited number of GW events currently available, we restrict our analysis to this effective approximation.
@@ 2.3 Inference of Local Merger Rate Density
Once the best-fitting DTD shape has been identified for each mass bin through the Bayesian framework described above, we proceed to infer the local merger rate density . For this inference, we note that the shape of determines the redshift evolution of the merger rate, while acts as a normalization parameter that scales the overall amplitude. With the best-fitting trajectory fixed, the inference of becomes a single-parameter problem for each mass bin. We adopt a flat prior on ranging from 1 to 100 Gpc-3 yr-1. The likelihood function follows the same Poisson formulation, with the expected counts now explicitly dependent on . This yields both point estimates and uncertainties for the local BBH merger rate density across different mass bins.
@@ 4 Results
In contrast, higher-mass binaries (-) exhibit peaked DTDs, with merger probabilities strongly concentrated at short to intermediate delays (2-6 Gyr) and a sharp suppression of long delays. The decay in the high for the high mass bin is coming from the fact that the merger rate is increasing as one moves from lower redshifts to higher redshift (see figure 2 red solid line.). Whereas at the lower delay time (for Gyrs), the constraints are weak (exhibiting a large error) due to non-detection of high redshift sources impacted by the selection function. In figure 2 we show the normalized redshift distributions of BBH mergers, comparing the observed GW catalog data (solid line) with the reconstructed trajectories (dashed line) derived from the best-evidence DTD inferred using our grid-based technique. The remarkable agreement between the two provides a strong self-consistency check on our framework, demonstrating that the inferred DTD not only yield statistically favored trajectories in space but also naturally reproduce the observed cosmic evolution of the merger population.